# Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions

### v2

@@ 3.  Safe Bidding Strategies
Recall that strategies chosen by the learner must be causal (mapping the history to a bidding strategy) while remaining RoI feasible for all rounds $t\in[T]$, even under adversarially generated competing bids. This creates an obvious challenge: a bidding strategy that satisfies RoI feasibility in previous rounds may become infeasible even under the slightest change in the competing bids e.g., if $\mathbf{v}=[0.9,0.5,0.1]$, the bidding strategy $(0.6,3)$ is RoI feasible for competing bids $\bm{\beta}_{-}^{1}=[0.61,0.59,0.59]$ but not for $\bm{\beta}_{-}^{2}=[0.59,0.59,0.59]$. To address this, we focus on safe bidding strategies that inherently guarantee RoI feasibility, regardless of the adversarial behavior of competing bids:
@@ 3.1.  Characterizing Safe Bidding Strategies
We begin by defining underbidding (overbidding) under the $m$-uniform bidding format in the given context. Recall that in a single-item auction, if the value of the item is $v$, then a bid $b$ is an underbid (overbid) if $b<v~{}(b>v)$.
@@
A $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an
(a) underbidding strategy if $b_{j}\leq w_{Q_{j}},\forall j\in[m]$ and $\exists\ell\in[m]$ such that $b_{\ell}<w_{Q_{\ell}}$.
(b) overbidding strategy if $\exists\ell\in[m]$ such that $b_{\ell}>w_{Q_{\ell}}$.
Figure 2: The solid line represents the average cumulative valuation curve, $\mathbf{w}$, and the dotted line represents the valuation curve, $\mathbf{v}$. The figure in the left (resp. right) illustrates underbidding (resp. overbidding) for a $2$-uniform bidding strategy. Note that the notions of underbidding and overbidding in Definition 3 are defined with respect to $\mathbf{w}$ and not $\mathbf{v}$ [^8] In Definition 3 (and throughout the paper), we assume $\gamma=0$ without loss of generality for ease of exposition. Footnote 8 illustrates how different values of $\gamma$ impacts the results. Specifically, for any $\gamma\geq 0$, underbidding and overbidding are defined with respect to the scaled average cumulative valuation curve, $\frac{\mathbf{w}}{1+\gamma}$.Here, the plots of $\mathbf{v}$ and $\mathbf{w}$ are shown to be linear for illustrative purposes only.
In other words, if $\gamma=0$, $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an underbidding strategy if, for all $j\in[m]$, the bid $b_{j}$ is at most the average of the first $Q_{j}$ entries of the valuation vector, denoted as $w_{Q_{j}}$, and there exists $\ell\in[m]$ where the inequality is strict. Recall that $Q_{j}=\sum_{\ell\leq j}q_{\ell}$ denotes the maximum number of demanded units in the first $j$ bid-quantity pairs. Similarly, $\mathbf{b}$ is an overbidding strategy if there exists some $\ell\in[m]$ such that $b_{\ell}$ is strictly greater than the average of the first $Q_{\ell}$ entries of the valuation vector. Having defined the notions of overbidding and underbidding, we characterize the class of safe bidding strategies, $\mathscr{S}_{m}$:
@@ A $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an
The proofs of this section are presented in Appendix A. Since there are infinitely many safe strategies in $\mathscr{S}_{m}$, we focus on the subset of safe undominated strategies, $\mathscr{S}_{m}^{\star}$, obtained by removing very weakly dominated strategies. A safe strategy $\mathbf{b}$ is said to be very weakly dominated if there exists another safe strategy $\mathbf{b}^{\prime}$ such that, for any competing bid $\bm{\beta}_{-}$, $V(\mathbf{b},\bm{\beta}_{-})\leq V(\mathbf{b}^{\prime},\bm{\beta}_{-})$ (Shoham and Leyton-Brown, 2008, pp. 79). Keeping these strategies does not improve the bidder’s performance, but removing them yields a finite strategy class (as shown below), which significantly simplifies the design of online learning algorithms. We now present the main result of this section:
@@
Observe that $\mathscr{S}_{m}^{\star}$ is a finite class of bidding strategies that depend only the valuation curve and is independent of the competing bids. The strategies in $\mathscr{S}_{m}^{\star}$ exhibit a “nested” structure, referred to as nested $m$-uniform bidding strategies hereafter, where the $j^{th}$ highest bid ($b_{j}$) is the average of the first $Q_{j}=\sum_{\ell=1}^{j}q_{\ell}$ entries of the valuation curve (illustrated in Fig. 1). Importantly, fixing $Q_{j}$’s uniquely determines the bidding strategy, as $b_{j}=w_{Q_{j}}$ for strategies in $\mathscr{S}_{m}^{\star}$ which will be crucial in learning the optimal safe bidding strategy, as discussed in the following section.
@@ 4.  Learning Safe Bidding Strategies
Here, the clairvoyant and the learner choose strategies from the class of safe bidding strategies with at most $m$ bid-quantity pairs, $\mathscr{U}_{m}^{\star}$. Thus, the regret defined in Eq. 3 is
$\mathsf{REG}=\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,.$
To design learning algorithms, we first consider the offline setting, where we aim to solve the following problem given the bid history, $[\bm{\beta}_{-}^{t}]_{t\in[T]}$.
@@
Since $\mathscr{U}_{m}^{\star}$ contains $O(M^{m})$ strategies, evaluating each strategy individually is intractable. To overcome this, we reduce the offline problem to finding the maximum weight path in a carefully constructed edge-weighted directed acyclic graph (DAG), which allows for the computation of the optimal offline bidding strategy in poly($m,M$) time.
4.1.  Offline Problem
Fix a bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, and consider (Offline). The following lemma shows that the objective function can be decomposed across the bid-quantity pairs. Formally,
@@ 4.1.  Offline Problem
The proofs of this section are presented in Appendix B. Building on the decomposition in Lemma 4.1, we now construct an edge-weighted DAG $\mathcal{G}(V,E)$ as proposed earlier.
Vertices. The DAG has a ‘layered’ structure with a source ($s$), destination ($d$) and $m$ intermediate layers each with $M$ nodes, denoted by $(\ell,j)$ where $\ell\in[m],j\in[M]$ as shown in Fig. 3. For convenience, we set $s=(0,0)$.
@@
$\mathsf{w}(e)=\sum_{t=1}^{T}\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$ (4)
(ii) $x=(\ell,j)$ and $y=d$, $\forall\ell\in[m],j\in[M]$. This edge connects the current layer to the destination node $d$, and its weight is $\mathsf{w}(e)=0$.
@@
  • There exists a bijective mapping between the $s$-$d$ paths in $\mathcal{G}(V,E)$ and bidding strategies in $\mathscr{U}_{m}^{\star}$, i.e., the path $s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for $k\in[m]$ refers to the strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
@@
    where $w_{z_{\ell}}$ is defined as per Eq. 1, $\forall\ell\in[k]$. Conversely, the strategy $\mathbf{b}=\langle(w_{Q_{1}},q_{1}),\dots,(w_{Q_{k}},q_{k})\rangle$ maps to the path $s\to(1,Q_{1})\to\dots\to(k,Q_{k})\to d$.
@@ 4.2.  Online Setting
In the online setting, in each round $t\in[T]$, the learner submits a strategy $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$ and receives feedback from the auction. We consider two feedback models: (a) full information setting and (b) bandit setting, with the latter discussed in Section B.3.1. Leveraging the DAG formulation, we design an online learning algorithm such that $\mathsf{REG}=o(T)$. Without loss of generaity, assume $v_{j}\in[0,1],\forall j\in[M]$.
In the full information setting, after receiving feedback at the end of round $t$, a DAG $\mathcal{G}^{t}(V,E)$ is constructed as described earlier. The edge weights are set as follows:
@@
$\mathsf{w}^{t}(e)=\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$ (5)
(ii) If $x=(\ell,j)$ and $y=d$, then $\mathsf{w}^{t}(e)=0,\forall\ell\in[m],j\in[M]$.
The online algorithm in Algorithm 1 consists of three main steps: (1) $\mathsf{UPDATE}$, (2) $\mathsf{SAMPLE}$, and (3) $\mathsf{MAP}$. In the first two steps, the algorithm maintains and updates the probabilities over the edges in the constructed DAG and selects a $s$-$d$ path in the DAG by sampling from these probabilities. These steps implement exponential weight updates using the weight-pushing method proposed by Takimoto and Warmuth (2003). Naively implementing exponential weight updates by treating each $s$-$d$ path (safe strategy) as an expert is intractable as there are $O(M^{m})$ such experts. However, following the decomposition in Lemma 4.1, Algorithm 1 only needs to maintain weights for individual edges and then sample a path based on these weights, avoiding the need to track weights for each path separately. The $\mathsf{MAP}$ step is the key component of the algorithm that maps the sampled path to a safe strategy (by Theorem 4.2).
Algorithm 1 Compute Bidding Strategies (Full Information)
0:  valuation curve $\mathbf{v}$, time horizon $T$, learning rate $\eta>0$.
1:  for $t=1,2,\dots,T$ do
2:     Construct $\mathcal{G}^{t}(V,E)$ similar to $\mathcal{G}(V,E)$ without weights.
3:     $\mathsf{UPDATE}:$ Obtain edge probabilities $\varphi^{t}(\cdot)$ by Algorithm 2. $\mathsf{SAMPLE}$: Define initial node $u=s$ and path $\mathfrak{p}=s$.
4:     while $u\neq d$ do
5:        Sample $v$ with probability $\varphi^{t}(u\to v)$.
6:        Append $v$ to the path $\mathfrak{p}$; set $u\leftarrow v$.
7:     end while
8:     $\mathsf{MAP}$: If $\mathfrak{p}=s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for some $k\in[m]$, submit $\mathbf{b}^{t}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
$b_{\ell}=w_{z_{\ell}}\quad\text{and}\quad q_{\ell}=z_{\ell}-z_{\ell-1},~{}\forall\ell\in[k]\,.$
where $w_{z_{\ell}}$ is defined as per Eq. 1 for all $\ell\in[k]$.
9:     The bidder observes $\bm{\beta}_{-}^{t}$ and sets $\mathsf{w}^{t}(e)$ as per Eq. 5.
10:  end for
@@
Algorithm 1 runs in $\text{poly}(m,M)$ time per round and achieves $\mathsf{REG}\leq O(M\sqrt{mT\log M})$ under full information feedback, and $\mathsf{REG}\leq O(M\sqrt{m^{3}MT\log M})$ in the bandit setting.
Algorithm 2 Update Edge Probabilities
0:  For $t\geq 1$, $\mathcal{G}^{t}(V,E)$. For $t\geq 2$, additionally require $\eta$, $\Gamma^{t-1}(v),\varphi^{t-1}(e),$ and $\mathsf{w}^{t-1}(e)$, $\forall e\in E$ and $\forall v\in V$.
1:  If $t=1$, – Initialize the edge probability of each edge $e=x\to y$ as $\varphi^{1}(e)=\nicefrac{{1}}{{|\mathcal{N}_{x}|}}$ where $\mathcal{N}_{x}$ is the set of out-neighbors of $x$.
2:  If $t\geq 2$: – Set $\Gamma^{t-1}(d)=1$ and recursively compute in bottom-to-top fashion for every node $u$ in $\mathcal{G}^{t}(V,E)$:
$\Gamma^{t-1}(u)=\sum_{v:u\to v=e\ni E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))$
– For edge $e=u\to v$ in $\mathcal{G}^{t}(V,E)$, update edge probability: $\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)}.$
We prove Theorem 4.3 by showing that Algorithm 1 is an efficient (polynomial time) and equivalent implementation of the exponential weight updates algorithm. We complement Theorem 4.3 by an establishing matching lower bound.
@@
Suppose $M\geq 2$ and $m=1$. Then, there exist competing bids, $[\bm{\beta}_{-}^{t}]_{t\in[T]}$, such that, under any learning algorithm, the expected regret $\mathbb{E}[\mathsf{REG}]=\Omega(M\sqrt{T})$ in both full information and bandit settings.

### v4

@@ 3. Safe Bidding Strategies
Recall that strategies chosen by the learner must be non-anticipating (mapping the history to a bidding strategy) while remaining RoI feasible for all rounds $t\in[T]$, even under adversarially generated competing bids. This creates an obvious challenge: a bidding strategy that satisfies RoI feasibility in previous rounds may become infeasible even under the slightest change in the competing bids e.g., if $\mathbf{v}=[0.9,0.5,0.1]$, the bidding strategy $(0.6,3)$ is RoI feasible for competing bids $\bm{\beta}_{-}^{1}=[0.61,0.59,0.59]$ but not for $\bm{\beta}_{-}^{2}=[0.59,0.59,0.59]$. To address this, we focus on safe bidding strategies that inherently guarantee RoI feasibility, regardless of the adversarial behavior of competing bids:
@@
Recall that we assume $\gamma=0$ throughout; otherwise the constraint in Definition 2 is identical to Eq. 2.
@@ 3.1. Characterizing Safe Bidding Strategies
We begin by defining underbidding (overbidding) under the $m$-uniform bidding format in the given context.
@@
A $m$-uniform strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an
(a) underbidding strategy if $b_{j}\leq w_{Q_{j}},\forall j\in[m]$ and $\exists\ell\in[m]$ such that $b_{\ell}<w_{Q_{\ell}}$, and
(b) overbidding strategy if $\exists\ell\in[m]$ such that $b_{\ell}>w_{Q_{\ell}}$,
where we recall that $Q_{j}=\sum_{\ell\leq j}q_{\ell},\forall j\in[m]$.
Figure 2: The solid line represents the average cumulative valuation curve, $\mathbf{w}$, and the dotted line represents the valuation curve, $\mathbf{v}$. The figure in the left (resp. right) illustrates underbidding (resp. overbidding) for a $2$-uniform bidding strategy. Note that the notions of underbidding and overbidding in Definition 3 are defined with respect to $\mathbf{w}$ and not $\mathbf{v}$. [^16] In Definition 3 (and throughout the paper), we assume $\gamma=0$ without loss of generality for ease of exposition. Footnote 16 illustrates how different values of $\gamma$ impacts the results. Specifically, for any $\gamma\geq 0$, underbidding and overbidding are defined with respect to the scaled average cumulative valuation curve, $\frac{\mathbf{w}}{1+\gamma}$. Here, the plots of $\mathbf{v}$ and $\mathbf{w}$ are shown to be linear for illustrative purposes only.
In other words, after normalizing $\gamma=0$, $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an underbidding strategy if, for all $j\in[m]$, the bid $b_{j}$ is at most the average of the first $Q_{j}$ entries of the valuation vector, denoted as $w_{Q_{j}}$, and there exists $\ell\in[m]$ where the inequality is strict. Similarly, $\mathbf{b}$ is an overbidding strategy if there exists some $\ell\in[m]$ such that $b_{\ell}$ is strictly greater than the average of the first $Q_{\ell}$ entries of the valuation vector. Having defined the notions of overbidding and underbidding, we characterize the class of safe bidding strategies, $\mathscr{S}_{m}$:
@@ A $m$-uniform strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an
Since there are infinitely many safe strategies in $\mathscr{S}_{m}$, we focus on the subset of safe undominated strategies, $\mathscr{S}_{m}^{\star}$, obtained by eliminating very weakly dominated strategies. A safe strategy $\mathbf{b}$ is said to be very weakly dominated if there exists another safe strategy $\mathbf{b}^{\prime}$ such that, for any competing bid $\bm{\beta}_{-}$, $V(\mathbf{b},\bm{\beta}_{-})\leq V(\mathbf{b}^{\prime},\bm{\beta}_{-})$ (Shoham and Leyton-Brown, 2008, pp. 79). Keeping these strategies does not improve the bidder’s performance, but removing them yields a finite strategy class (as shown below), which significantly simplifies the design of online learning algorithms. We show that for every underbidding strategy (as per Definition 3) in $\mathscr{S}_{m}$, there exists a non-underbidding strategy in the same class that very weakly dominates it, allowing the former to be safely eliminated. Hence,
@@
Observe that $\mathscr{S}_{m}^{\star}$ is a finite class of bidding strategies that depend only the valuation curve and is independent of the competing bids. The strategies in $\mathscr{S}_{m}^{\star}$ exhibit a “nested” structure, where the $j^{th}$ highest bid ($b_{j}$) is the average of the first $Q_{j}=\sum_{\ell=1}^{j}q_{\ell}$ entries of the valuation curve (illustrated in Fig. 1). Importantly, fixing $Q_{j}$’s uniquely determines the bidding strategy, as $b_{j}=w_{Q_{j}}$ for strategies in $\mathscr{S}_{m}^{\star}$, which will be crucial in learning the optimal safe bidding strategy, as discussed in the following section.
@@ 4. Learning Safe Bidding Strategies
Here, the clairvoyant and the learner choose strategies from the class of safe bidding strategies with at most $m$ bid-quantity pairs, $\mathscr{U}_{m}^{\star}$ (defined in Theorem 3.2). Thus, in this case, the regret defined in Eq. 3 becomes
$\textsf{REG}=\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,.$ (4)
To design learning algorithms, we first consider the offline setting, where we aim to solve the following problem for a given bid history, $\bm{\beta}_{-}^{t}]_{t\in[T]}$.
@@
Since $\mathscr{U}_{m}^{\star}$ contains $O(M^{m})$ strategies, evaluating each strategy individually is intractable. To overcome this, we reduce the offline problem to finding the maximum weight path in an appropriate edge-weighted directed acyclic graph (DAG), which allows for the computation of the optimal offline strategy in poly($m,M$) time.
4.1. Offline Setting
Fix a bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, and consider the Problem (Offline). The following lemma shows that the objective function can be decomposed across the bid-quantity pairs. Formally,
@@ 4.1. Offline Setting
Proof.
By Theorem 3.2, a bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell})\rangle\in\mathscr{U}_{m}^{\star}$ for some $\ell\in[m]$ can be uniquely identified by $\{Q_{1},\dots,Q_{\ell}\}$ where $Q_{j}=\sum_{k\leq j}q_{k}$ because $b_{j}=w_{Q_{j}}$ for all $j\in[\ell]$. Let the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$. Observe that if the bidder wins $r$ units in any auction, then the $r^{th}$ highest bid value in its submitted strategy must be at least the $r^{th}$ smallest bid in the top $K$ competing bids, i.e., $\bm{\beta}_{-,t}^{-(r)}$. Hence, for any round $t$,
$V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{j=1}^{\ell}\sum_{k=Q_{j-1}+1}^{Q_{j}}v_{k}\cdot\mathbb{I}{\left[w_{Q_{j}}\geq\bm{\beta}_{-,t}^{-(k)}\right]},$
which implies that the problem (Offline) can be equivalently expressed as
Building on the decomposition in Lemma 4.1, we now construct an edge-weighted DAG $\mathcal{G}(V,E)$.
Vertices. The DAG has a ‘layered’ structure with a source node ($s$), destination node ($d$) and $m$ intermediate layers. The $\ell^{th}$ intermediate layer contains $M+1$-$\ell$ nodes, denoted by $(\ell,j)$ where $\ell\in[m],j\in\{\ell,\ell+1,\dots,M\}$ as shown in Fig. 3. For convenience, we set $s=(0,0)$.
@@
$\textsf{w}(e)=\sum_{t=1}^{T}\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$ (5)
(ii) $x=(\ell,j)$ and $y=d$, $\forall\ell\in[m],j\in[M]$. This edge connects the current layer to the destination node $d$, and its weight is $\textsf{w}(e)=0$.
@@
There exists a bijective mapping between the $s$-$d$ paths in $\mathcal{G}(V,E)$ and strategies in $\mathscr{U}_{m}^{\star}$, i.e., the path $s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for $k\in[m]$ refers to the strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
@@
where $w_{z_{\ell}}$ is defined as per Eq. 1 and $z_{0}=0$. Conversely, the strategy $\mathbf{b}=\langle(w_{Q_{1}},q_{1}),\dots,(w_{Q_{k}},q_{k})\rangle$ maps to the path $s\to(1,Q_{1})\to\dots\to(k,Q_{k})\to d$.
@@ 4.2. Online Setting
In the online setting, in each round $t\in[T]$, the learner submits a strategy $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$ and receives feedback from the auction. We consider two feedback models: (a) full information setting and (b) bandit setting. Leveraging the DAG formulation, we design a polynomial-time online learning algorithm such that $\textsf{REG}=\widetilde{O}(\text{poly}(m,M)\cdot\sqrt{T})$. For the rest of the work, we assume $v_{j}\in[0,1],\forall j\in[M],$ without loss of generality.
In Theorem 4.2, we established a bijection mapping between ssdd paths in $\mathcal{G}(V,E)$ and safe strategies in $\mathscr{U}_{m}^{\star}$. While naïvely applying the Hedge algorithm (Freund and Schapire, 1997) in the full-information setting by treating each path as an expert yields optimal regret bounds, it is computationally intractable due to the exponential number of paths: $O(M^{m})$. The situation further worsens in the bandit setting, as implementing the EXP3 algorithm (Auer et al., 2002), treating each path as an expert, also yields unsatisfactory regret bounds (recall that EXP3 obtains $O(\sqrt{NT\log N})$ regret where $N$ is the number of arms and per-round rewards belong to $[0,1]$. In this setting, $N=O(M^{m})$.).
Our proposed algorithm retains the core idea of treating paths as experts but does so efficiently. Leveraging the decomposition in Lemma 4.1, it maintains probabilities over edges rather than paths, enabling efficient path sampling as a Markov chain without explicitly maintaining a distribution over all the paths. Specifically, in every round $t$, a DAG $\mathcal{G}^{t}(V,E)$, as described earlier, is constructed. The proposed algorithm consists of three main steps: (1) UPDATE, (2) SAMPLE, and (3) MAP. In the UPDATE step, the bidder uses the feedback from the previous round to update the edge probabilities, $\varphi^{t}(u\to v)$, for every edge $e=u\to v$ in the DAG. Using the dynamic programming-based weight-pushing method from Takimoto and Warmuth (2003), this step can be performed in $\text{poly}(m,M)$ time. In the SAMPLE step, the algorithm samples an ssdd path in a Markovian manner: starting at the source node $s$, it repeatedly samples the next node $v$ from the set of neighbors of the current node $u$ according to probabilities $\varphi^{t}(u\to v)$, until the destination node $d$ is reached. Once a $s$-$d$ path is sampled, in the MAP step, the bidder maps it to a safe strategy $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$ and submits $\mathbf{b}^{t}$ in round $t$. After receiving feedback at the end of the round, in the full information setting, the edge weights are set as follows:
@@
$\textsf{w}^{t}(e)=\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$ (6)
(ii) If $x=(\ell,j)$ and $y=d$, then $\textsf{w}^{t}(e)=0,\forall\ell\in[m],j\in[M]$.
The proposed algorithm is formally presented in Algorithm 1. The main result of this section is:
Algorithm 1 Learning Safe Bidding Strategies (Full Information)
1:valuation curve $\mathbf{v}$, time horizon $T$, learning rate $\eta\in(0,\frac{1}{M}]$, $\textsf{w}^{0}(e)=0,\text{and}~\varphi^{0}(e)=1,\forall e\in E$.
2:for $t=1,2,\dots,T$ do
3:  Construct $\mathcal{G}^{t}(V,E)$ similar to $\mathcal{G}(V,E)$ without weights.
4:  $\textsf{UPDATE}:$ Set $\Gamma^{t-1}(d)=1$ and recursively compute in bottom-to-top fashion for every node $u$ in $\mathcal{G}^{t}(V,E)$:
$\Gamma^{t-1}(u)=\sum_{v:u\to v=e\ni E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))\,.$
5:  For edge $e=u\to v$ in $\mathcal{G}^{t}(V,E)$, update edge probability:
$\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)}.$
6:  SAMPLE: Define initial node $u=s$ and path $\mathfrak{p}^{t}=s$.
7:  while $u\neq d$ do
8:    Sample $v$ with probability $\varphi^{t}(u\to v)$.
9:    Append $v$ to the path $\mathfrak{p}^{t}$; set $u\leftarrow v$.
10:  end while
11:  MAP: If $\mathfrak{p}^{t}=s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for some $k\in[m]$, submit $\mathbf{b}^{t}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
$b_{\ell}=w_{z_{\ell}}\qquad\text{and}\qquad q_{\ell}=z_{\ell}-z_{\ell-1},~\forall\ell\in[k]\,.$
where $w_{z_{\ell}}$ is defined as per Eq. 1 for all $\ell\in[k]$ and $z_{0}=0$.
12:  The bidder observes $\bm{\beta}_{-}^{t}$ and sets $\textsf{w}^{t}(e)$ as per Eq. 6.
13:end for
@@
In the full information setting, Algorithm 1 achieves $\textsf{REG}\leq O(M\sqrt{mT\log M})$ in $\text{poly}(m,M)$ space and time per round.
In the bandit setting, as the bidder receives only $x(\bm{\beta}^{t})$ and $P(\bm{\beta}^{t})$ as feedback in round $t$, they can not compute Eq. 6 for all the edges. Hence, we use an unbiased estimator, $\widehat{\textsf{w}}^{t}(e)$, of $\textsf{w}^{t}(e)$ as follows:
$\widehat{\textsf{w}}^{t}(e)$ $=\overline{\textsf{w}}(e)-\frac{\overline{\textsf{w}}(e)-\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}\quad\text{where}\quad p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\,.$ (7)
Here, $p^{t}(e)$ is the probability of selecting edge $e$ in round $t$, $\mathbb{P}^{t}$ is the distribution over all $s$-$d$ paths in $\mathcal{G}^{t}(V,E)$. Observe that $\widehat{\textsf{w}}^{t}(e)$ is well defined as $p^{t}(e)>0,\forall e\in E,\forall t\in[T]$. Here, $\overline{\textsf{w}}(e)$ is an edge-dependent upper bound of the edge weight $\textsf{w}^{t}(e)$. Formally, for $\ell\in[m]$,
$\overline{\textsf{w}}(e)=\begin{cases}j^{\prime}-j,&\text{if }e=(\ell-1,j)\to(\ell,j^{\prime})\\ 0,&\text{if }e=(\ell,j)\to d\,.\end{cases}$
The algorithm followed by the bidder in the bandit setting is identical to Algorithm 1, with each instance of $\textsf{w}^{t}(e)$ replaced by $\widehat{\textsf{w}}^{t}(e),\forall t\in[T],\forall e\in E$. Furthermore, define $\widehat{\textsf{w}}^{0}(e)=0,\forall e\in E$. Clearly, $\mathbb{E}[\widehat{\textsf{w}}^{t}(e)]=\textsf{w}^{t}(e)$. The unbiased estimator $\widehat{\textsf{w}}^{t}(e)$ is different from the natural importance-weighted estimator $\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}$ for technical reasons (similar to Lattimore and Szepesvári (2020, Eq. 11.6)) which we will clarify in the regret analysis (see details in Section B.3). Hence,
@@
In the bandit setting, Algorithm 1 with the aforementioned changes achieves $\textsf{REG}\leq O(M^{2}\sqrt{m^{3}T\log M})$ in $\text{poly}(m,M)$ space and time per round.
Note that the per-round reward $V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})\in[0,M]$ and number of experts is $N=O(M^{m})$. In the bandit setting, although we compute unbiased estimators of the edge weights (equivalently path weights) and perform Hedge-like updates efficiently, the resulting regret bound is tighter than that of a naïve EXP3 implementation. This improvement stems from two key advantages: (a) after each round, the bidder can observe the rewards of the constituent edges of the submitted path, rather than just the aggregate reward, [^17] Strictly speaking, this corresponds to the semi-bandit feedback model in online combinatorial optimization (Neu and Bartók, 2013; Jourdan et al., 2021; Brandt et al., 2022), whereas the full bandit model reveals only the total reward. However, in our setting, the full bandit model lacks a natural interpretation. For simplicity, we refer to this partial feedback setting as the bandit feedback model. and (b) the structure of the combinatorial action space reveals partial information about unselected paths as well. We complement Theorem 4.3 and Theorem 4.4 by establishing the following lower bound:
Theorem 4.5.
Suppose $M\geq 2$ and $m=1$. Then, there exist competing bids, $\bm{\beta}_{-}^{t}]_{t\in[T]}$, such that, under any learning algorithm, $\mathbb{E}[\textsf{REG}]=\Omega(M\sqrt{T})$ in the full information setting. This implies that the lower bound also holds under bandit feedback.